# Gauge dependence of momentum running in higher-derivative gravity

### v1

@@ IV Analysis
The gauge dependence observed in the last section disappears if we go on-shell with the background metric configuration. Let us consider the variation of the effective action with respect to the gauge fixing and the Lautrup-Nakanishi operators [30]:
@@
$\delta_{\xi}g_{\mu\nu}=\textbf{R}^{\mu\nu}{}_{\alpha}\xi^{\alpha}=\nabla^{(\nu}g^{\mu)}{}_{\alpha}\xi^{\alpha}\ ,$ (45)
[…]
@@
therefore only the $\mu$-running of the EH term and the boundary term $\square R$, neglected in our computation, can be gauge dependent [30, 33].
@@
In Yang-Mills theories and in the UV divergent part of quadratic gravity and conformal gravity, the background field method [8] permits one to extract from it the running of the couplings governing, for example, the 4-point amplitude, the simplest on-shell non-trivial process. This happens because the resulting theory for the background field is invariant with respect to background gauge transformations, and the UV divergences have to be strictly local. These two constraints, together with dimensional analysis, fix the structure of the YM background one-loop action to only $F^{2}$, the UV divergent part of the quadratic gravity action to the form $R^{2}+C^{2}$, and that of conformal gravity to only $C^{2}$. The UV divergent parts of both the 2-point function and the 4-point amplitude come from these operators, hence, they must be governed by the same coupling.
@@
If $n=1,2$, they contribute at one-loop to the 4-point on-shell scattering, but not to the background propagator. In this case, the resulting on-shell scattering amplitude is not proportional to the vertex generated by the local terms $R^{2}+C^{2}$ (or only $C^{2}$ in conformal gravity), and its dependence on momenta is not completely determined by their beta functions, even when taking into account IR contributions. The beta functions of the local operators, determined by the momentum dependence of the 2-point functions, are no longer associated with an on-shell gauge-invariant process, and can show a gauge dependence, as explicitly observed above and in [6]. That means the beta functions of the local operators are not enough to establish the UV behaviour of the theory, contrary to what happens in YM theories. To have something gauge invariant, we must compute on-shell one-loop scattering amplitudes, as done in [6] for the scalar sector of quadratic gravity.
@@ V Conclusions
In accordance with [6], we have found that the momentum-induced beta functions, which some of us claimed to be physical in previous works [1, 2], carry explicit dependence on unphysical parameters such as the $g_{i}$’s. This gauge dependence is due to the presence of logarithmically enhanced nonlocal operators in the one-loop effective action. They contribute to the on-shell $n$-point scattering amplitudes with $n>2$, which are the truly physical objects, and mix with local ones. Only the resulting combination of contributions needs to be gauge independent, not the single beta functions taken one by one.
@@
In [6], the 4-point amplitudes for the scalar mode were computed and were found to be also parametrization dependent. This, together with the presence of infrared divergences, could be the signal that the canonical degrees of freedom of quadratic gravity are not the correct asymptotic states of the theory in the high-energy regime. It could be that considering inclusive external states where ghost and massless particles are mixed, as suggested in [34], cancels the IR large logs from scattering amplitudes.

### v2

@@ IV Analysis
The gauge dependence observed in the last section disappears if we go on-shell with the background metric configuration. Let us consider the variation of the effective action with respect to the gauge fixing and the Lautrup-Nakanishi operators [Avramidi:1986mj]:
@@
$\delta_{\xi}g^{\mu\nu}=\textbf{R}^{\mu\nu}{}_{\alpha}\xi^{\alpha}=\nabla^{(\nu}g^{\mu)}{}_{\alpha}\xi^{\alpha}\ ,$ (45)
$\textbf{R}^{\mu\nu}{}_{\alpha,\rho\sigma}$ is its functional derivative with respect to the metric, and $F_{\alpha,\mu\nu}$ is the functional derivative of the gauge fixing $F_{\alpha}$ with respect to the fluctuation field $h_{\mu\nu}$. The equations of motion have mass dimension 4, because the metric is taken dimensionless. Since the expression in the first line of (44) can be reduced at the leading order to the inverse of a $4\text{th}$ order differential operator contracted with $\epsilon_{\gamma\delta}$, its UV divergence is given by a tadpole diagram and must be proportional to the trace of $\epsilon$ [Barvinsky:1985an, Avramidi:1986mj]. In the second line, the first term has the same type of divergence, while the second term is composed of two equations of motion multiplied by the inverse of a differential operator of rank $8$, so it is not UV divergent at all. That means the gauge dependence of UV divergences is proportional to the trace of the equations of motion.
@@
therefore only the $\mu$-running of the EH term and the boundary term $\square R$, neglected in our computation, can be gauge dependent [Avramidi:1986mj, Voronov:1984kq].
@@
In Yang-Mills theories and in the UV divergent part of quadratic gravity and conformal gravity, the background field method [Weinberg:1996kr] permits one to extract from it the running of the couplings governing, for example, the 4-point amplitude, the simplest on-shell non-trivial process. This happens because the resulting theory for the background field is invariant with respect to background gauge transformations, and the UV divergences have to be strictly local. These two constraints, together with dimensional analysis, fix the structure of the YM background one-loop action to only $F^{2}$, the UV divergent part of the quadratic gravity action to the form $R^{2}+C^{2}$, and that of conformal gravity to only $C^{2}$. The UV divergent parts of both the 2-point function and the 4-point amplitude come from these operators, hence, they must be governed by the same coupling.
@@
If $n=1,2$, they contribute at one-loop to the 4-point on-shell scattering, but not to the background propagator. In this case, the resulting on-shell scattering amplitude is not proportional to the vertex generated by the local terms $R^{2}+C^{2}$ (or only $C^{2}$ in conformal gravity), and its dependence on momenta is not completely determined by their beta functions, even when taking into account IR contributions. The beta functions of the local operators, determined by the momentum dependence of the 2-point functions, are no longer associated with an on-shell gauge-invariant process, and can show a gauge dependence, as explicitly observed above and in [Salvio:2025cmi]. That means the beta functions of the local operators are not enough to establish the UV behaviour of the theory, contrary to what happens in YM theories. To have something gauge invariant, we must compute on-shell one-loop scattering amplitudes, as done in [Salvio:2025cmi] for the scalar sector of quadratic gravity.
@@ V Conclusions
In accordance with [Salvio:2025cmi], we have found that the momentum-induced beta functions, which some of us claimed to be physical in previous works [Buccio:2024hys, Buccio:2024omv], carry explicit dependence on unphysical parameters such as the $g_{i}$’s. This gauge dependence is due to the presence of logarithmically enhanced nonlocal operators in the one-loop effective action. They contribute to the on-shell $n$-point scattering amplitudes with $n>2$, which are the truly physical objects, and mix with local ones. Only the resulting combination of contributions needs to be gauge independent, not the single beta functions taken one by one.
@@
In [Salvio:2025cmi], the 4-point amplitudes for the scalar mode were computed and were found to be also parametrization dependent. This, together with the presence of infrared divergences, could be the signal that the canonical degrees of freedom of quadratic gravity are not the correct asymptotic states of the theory in the high-energy regime. It could be that considering inclusive external states where ghost and massless particles are mixed, as suggested in [Holdom:2021hlo], cancels the IR large logs from scattering amplitudes.